### v1

@@ Abstract
Keywords:
6DOF object pose estimation Dataset creation Unsupervised learning
@@ 1 Introduction
Articulated objects, comprising multiple parts connected by revolute or prismatic joints with varying joint states (rotational angle of a revolute joint or translation length of a prismatic joint), commonly exist in the real world. The interactions between humans and these objects give rise to numerous practical applications, such as robot manipulations and automation in industrial processes [5, 25]. Therefore, pose estimation for these objects has become a crucial problem in computer vision. We focus on accomplishing the category-level articulated object pose estimation through a self-supervised approach. Our objective is to use a point cloud of unknown articulated objects within known categories obtained from a single-frame RGB-D image segmented by detection models such as Mask-RCNN [9] as input. Then, we infer each part’s pose and segmentation, each joint’s direction and pivot, as illustrated in Fig. 1. We aim to achieve this without utilizing pose and shape annotations during training. Due to the varying shapes, poses, and complex real-world environments, this task is ill-posed and remains challenging.
Many works have focused on solving the aforementioned task under simpler problem settings. Unsupervised Pose-aware Part Decomposition (UPPD) [16] utilizes object shape annotations as a substitute for pose annotations. PartMobility [35] utilizes multiple-frame point clouds of the same object under different joint states. However, these methods still face limitations when confronted with scenarios where shape information is unavailable or when dealing with single-frame data. To the best of our knowledge, Equi-Articulated-Pose (EAP) [24] is the only work that has tackled this task with single-frame point cloud as input and without shape or pose annotations on a synthetic dataset. EAP guides the network to learn part-by-part reconstruction of the input shapes by combining disentangled information, such as canonical part shapes, object structure, and part-level poses, in a self-supervised manner. To achieve such disentanglement, EAP extracts part-level $\mathrm{SE}(3)$-equivariant shape feature of a local region, instead of object-level $\mathrm{SE}(3)$-equivariant one, from an input and part-level poses. Since part-level poses are not given in nature, EAP requires iterative updates of such poses. It also uses an inner iterative operation, Slot-Attention [26], for segmenting parts. These iterative operations sacrifice inference speed.
@@
PartMobility [35]
UPPD [16]
EAP [24]
@@ 2 Related Works
Category-level rigid object pose estimation: This task focuses on predicting an unknown rigid object’s pose from images. NOCS [39] predicts the per-pixel coordinates in canonical space from RGB-D images. Several methods [13, 12, 37] further employ CAD models from ShapeNet dataset [2] to generate shape templates and use iterative closest point (ICP) [34] for matching the pose. Commonly used backbone for this task is 3D graph convolution network (3DGCN) [22] and PointNet++ [33]. These methods require expensive large-scale dataset annotation. Some approaches attempt to accomplish this task in a self-supervised manner. With the CAD model available, several methods [36, 38] render the predicted pose with the CAD model as a synthetic image and compare it with the input image. Some methods focus on the multi-view RGB images provided cases [11, 19]. Especially, $\mathrm{SE}(3)$-eSCOPE [21] achieved this task with single-view input and without pose annotations or CAD models. They use the $\mathrm{SE}(3)$-equivariant backbone, Equivariant Point Net (EPN) [3], to simultaneously conduct $\mathrm{SE}(3)$-invariant shape reconstruction as a reference frame, and predict the $\mathrm{SE}(3)$-equivariant pose transformation which sends input to the reconstruction.
@@
Category-level articulated object pose estimation: This task focuses on predicting part-level pose, part-level segmentation, and joint information for unknown objects within known categories. Previous methods [20, 14, 1, 41] try to solve this task with RGB-D image or video input by directly estimating the part-level pose. Some methods [29, 18, 31] transfer the task into a movable shape reconstruction task with neural implicit representation [27, 30] and predict the pose indirectly. Some methods [8, 15] parameterize the joint movement with active interaction with articulated objects. To reduce the segmentation cost, [23] uses semantic segmentation annotation and transfers it into part segmentation to conduct semi-supervised learning. However, similar to the rigid object pose estimation, the cost of the dataset annotation limits the application of these methods. To solve this task in a self-supervised manner, UPPD [16] utilizes the annotation of object shape instead of the annotation of object pose. Some methods [35, 10] used multi-view observation with the same object in different joint states to predict the joint movement. EAP [24] solved such a task with a single-frame point cloud input and without shape or pose annotation. EAP repeats the process of segmenting each part, reconstructing the per-part $\mathrm{SE}(3)$-invariant shape, and predicting the per-part pose multiple times to gain a refined pose estimation. However, directly segmenting parts for inputs with different poses and shapes is challenging, often resulting in poor accuracy, and the inference speed is unsuitable for real-time applications.
@@
Articulated Object Dataset: Synthetic datasets of articulated objects such as Shape2Motion, SAPIEN, and PartNet [40, 43, 28] are commonly used in the articulated object pose estimation. Compared to RGB-D images captured from the real world, these datasets lack the consideration of complicated real-world environments. HOI4D [25] collects multiple articulated and rigid object mesh data and RGB-D images in human-object interaction. However, due to the mismatch between the depth and RGB channels, a non-negligible amount of noise is present in their ground-truth annotation of part segmentation based solely on the RGB channels.
@@ 3 Method
Figure 2: Pipeline of OP-Align. At the object-level phase, for the input point cloud $\mathbf{X}$, we use the E2PN [44] backbone to predict and select object-level pose $\mathbf{R}_{\mathrm{o}},\mathbf{t}_{\mathrm{o}}$ from pose candidates, and generate the canonical reconstruction $\mathbf{Y}$ by adding a learnable parameter called category-common base shape $\mathbf{Y}_{\mathrm{base}}$. At the part-level phase, two PointNets [32] with shared weights predict the part segmentation probability $\mathbf{W}_{\mathrm{x}},\mathbf{W}_{\mathrm{y}}$, joint states $\mathbf{a}_{\mathrm{x}},\mathbf{a}_{\mathrm{y}}$, joint pivots $\mathbf{c}_{\mathrm{x}},\mathbf{c}_{\mathrm{y}}$, and joint directions $\mathbf{d}_{\mathrm{x}},\mathbf{d}_{\mathrm{y}}$ for object-level aligned input $\mathbf{R}_{\mathrm{o}}\mathbf{X}+\mathbf{t}_{\mathrm{o}}$ and reconstruction $\mathbf{Y}$, to generate part-level alignment $\mathbf{R}_{\mathrm{d}},\mathbf{R}_{\mathrm{a}},\mathbf{T}_{\mathrm{a}}$ that aligns each part of $\mathbf{X}$ to the corresponding part of $\mathbf{Y}$ as part-level aligned inputs $\mathbf{Z}$.
Category-level articulated object pose estimation can be defined as follows. Given a point cloud $\mathbf{X}\in\mathbb{R}^{3\times N}$ of an articulated object consisting of $P$ parts, we assign each point to a part, predict the rotation and translation for each part, and provide the pivot and the direction for each joint. To solve this problem, our model predicts each point’s segmentation probability $\mathbf{W}\in\mathbb{R}^{P\times N}$, each joint’s pivot and direction $\mathbf{\{}\mathbf{c}_{[i]}\in\mathbb{R}^{3},\mathbf{d}_{[i]}\in\mathbb{R}^{3}\mid i\in\{1,2,\dots,J\}\}$, and the rotation and the translation for each part $\{\mathbf{R}_{[i]}\in\mathrm{SO}(3),\mathbf{t}_{[i]}\in\mathbb{R}^{3}\mid i\in\{1,2,\dots,P\}\}$. During training, we assume that the number of parts $P$ and the type of joints (revolute or prismatic) are given. Specifically, OP-Align assumes that each joint connects two independent parts, resulting in $J=P-1$ joints, which cover most of the articulated object categories found in daily environments.
The pipeline of OP-Align is shown in Fig. 2. At the object-level phase, OP-Align initially employs Efficient $\mathrm{SE}(3)$-equivariant Point Net (E2PN) [44] for object-level pose selection from a discretization of the $\mathrm{SE}(3)$ group, and generate canonical reconstruction. At the part-level phase, two PointNets (PNs) [32] with shared weights perform part segmentation and joint parameters estimation separately for the input aligned with object-level pose and the canonical reconstruction. The obtained joint parameters generate the part-level alignment between the input and the canonical reconstruction, aligning each part of the input with its corresponding part of the reconstruction by simulating the joint movement.
@@
Notice in this section, for a rank $n$ tensor $A$, we denote the $(i_{1},i_{2},\dots,i_{n})$-element (a rank 00-tensor) as $A_{[i_{1},i_{2},\dots,i_{n}]}$. Moreover, we use NumPy [7] like notation to extract a tensor from $A$ (but each index starts from $1$). For example, $A_{[i_{1}]}$ denote the $i_{1}$-th rank $(n-1)$ tensor along the first axis and $A_{[:,i_{2}]}$ denote the $i_{2}$-th rank $(n-1)$ tensor along the second axis.
@@ 3.1.1 Expansion from rigid objects to articulated objects
To solve the rigid object pose estimation in a self-supervised manner, $\mathrm{SE}(3)$-eSCOPE [21] utilizes a $\mathrm{SE}(3)$-equivariant backbone to disentangle shape and pose by generating an $\mathrm{SE}(3)$-invariant shape reconstruction and selecting $\mathrm{SE}(3)$-equivariant pose from candidates in a discretization of the $\mathrm{SE}(3)$ group for aligning the reconstruction and input. They observed that poses of $\mathrm{SE}(3)$-invariant reconstructions for objects in the same category are often consistently aligned. However, for articulated objects, each part’s pose is also influenced by joint movement. This complexity renders the reconstruction generated by the $\mathrm{SE}(3)$-eSCOPE unable to maintain consistent poses for all the parts.
@@ 3.1.2 Weighted Point Cloud Distances
We combine part segmentation probability with the point cloud distance between the part-level aligned inputs and the reconstruction to learn part segmentation and part alignment simultaneously. To achieve this, we use weighted point cloud distances, and later, part segmentation probability will sometimes be set as weights. A commonly used point cloud distance is the chamfer distance (CD), and we also employ the Density-awarded Chamfer Distance (DCD) [42]. Given two point clouds $\mathbf{P}$ and $\mathbf{Q}$, the single-directional weighted CD (L1) and DCD from $\mathbf{P}$ to $\mathbf{Q}$ with the weight $\mathbf{w}$ are defined as
$\begin{split}\mathrm{CD}(\mathbf{P},\mathbf{Q},\mathbf{w})=&\frac{1}{|\mathbf{P}|}\sum_{n=1}^{|\mathbf{P}|}\mathbf{w}_{[n]}\min_{m\in\{1,2,\dots,|\mathbf{Q}|\}}\left\|\mathbf{P}_{[n]}-\mathbf{Q}_{[m]}\right\|,\\ \mathrm{DCD}(\mathbf{P},\mathbf{Q},\mathbf{w},\alpha)=&\frac{1}{|\mathbf{P}|}\sum_{n=1}^{|\mathbf{P}|}\mathbf{w}_{[n]}\min_{m\in\{1,2,\dots,|\mathbf{Q}|\}}\left(1-e^{-\alpha\left\|\mathbf{P}_{[n]}-\mathbf{Q}_{[m]}\right\|_{2}}\right).\\ \end{split}$ (1)
@@ 3.2 Object-level phase
In the Object-level phase, OP-Align performs object-level pose selection, following a methodology similar to $\mathrm{SE}(3)$-eSCOPE [21], and generate canonical reconstruction.
By feeding the input $\mathbf{X}$, E2PN [44] backbone initially outputs the $\mathrm{SE}(3)$-equivariant feature $\mathbf{F}_{\mathrm{eqv}}\in\mathbb{R}^{D\times 60}$. This feature is generated by $60$ anchors representing different poses of the object. Here, $60$ is the number of elements of the icosahedral rotation group, a discretization of the $3$D rotation group $\mathrm{SO}(3)$. Then, we max pool $\mathbf{F}_{\mathrm{eqv}}$ among anchor dimension to obtain a $\mathrm{SE}(3)$-invariant feature $\mathbf{f}_{\mathrm{inv}}\in\mathbb{R}^{D}$. We use PoseHead, consisting of multi-layer perceptron (MLP), to output per-anchor rotation and translation $\{(\mathbf{R}_{[i]},\mathbf{t}_{[i]})=\textrm{PoseHead}(\mathbf{F}_{\mathrm{eqv}[:,i]})\mid i\in\{1,2,\dots,60\}\}$. To obtain the canonical reconstruction $\mathbf{Y}\in\mathbb{R}^{3\times N}$, we also use an MLP called ReconHead and a learnable parameter $\mathbf{Y}_{\mathrm{base}}$ which represents the category-common base shape and is of the same size as $\mathbf{Y}$. The canonical reconstruction $\mathbf{Y}$ is obtained by adding the output of ReconHead and $\mathbf{Y}_{\mathrm{base}}$; $\mathbf{Y}=\textrm{ReconHead}(\mathbf{f}_{\mathrm{inv}})+\mathbf{Y}_{\mathrm{base}}$.
@@ 3.3 Part-level phase
OP-Align uses two PNs [32] with shared weights to process the object-level aligned input $\mathbf{R}_{\mathrm{o}}\mathbf{X}+\mathbf{t}_{\mathrm{o}}$ and the reconstruction $\mathbf{Y}$ separately. These two PNs output the segmentation probabilities $\mathbf{W}_{\mathrm{x}},\mathbf{W}_{\mathrm{y}}\in\mathbb{R}^{P\times N}$, joint pivots $\mathbf{c}_{\mathrm{x}},\mathbf{c}_{\mathrm{y}}\in\mathbb{R}^{(P-1)\times 3}$, joint directions $\mathbf{d}_{\mathrm{x}},\mathbf{d}_{\mathrm{y}}\in\mathbb{R}^{(P-1)\times 3}$ and per-part joint states $\mathbf{a}_{\mathrm{x}},\mathbf{a}_{\mathrm{y}}\in\mathbb{R}^{(P-1)\times 2}$ from each joint, where subscripts $\mathrm{x}$ and $\mathrm{y}$ indicate outputs from $\mathbf{R}_{\mathrm{o}}\mathbf{X}+\mathbf{t}_{\mathrm{o}}$ and $\mathbf{Y}$ respectively. Here, joint state $\mathbf{a}_{*}$ represents joint angles for revolute joints and translation lengths for prismatic joints, and the dimension of the second axis of $\mathbb{R}^{(P-1)\times 2}$ reflects the assumption that each joint connect two parts. We define the part-level aligned inputs $\mathbf{Z}_{[j,i]}$, $j=1,2,\dots,P-1$, $i=1,2$, obtained by a relative transformation that aligns the $i$-th part connected to the $j$-th joint of $\mathbf{R}_{\mathrm{o}}\mathbf{X}+\mathbf{t}_{\mathrm{o}}$ with the corresponding part of $\mathbf{Y}$ by

### v2

@@ 1 Introduction
Articulated objects, comprising multiple parts connected by revolute or prismatic joints with varying joint states (rotational angle of a revolute joint or translation length of a prismatic joint), commonly exist in the real world. The interactions between humans and these objects give rise to numerous practical applications, such as robot manipulations and automation in industrial processes [command, hoi4d]. Therefore, pose estimation for these objects has become a crucial problem in computer vision. We focus on accomplishing the category-level articulated object pose estimation through a self-supervised approach. Our objective is to use a point cloud of unknown articulated objects within known categories obtained from a single-frame RGB-D image segmented by detection models such as Mask-RCNN [MaskRCNN] as input. Then, we infer each part’s pose and segmentation, each joint’s direction and pivot, as illustrated in Fig. 1. We aim to achieve this without utilizing pose and shape annotations during training. Due to the varying shapes, poses, and complex real-world environments, this task is ill-posed and remains challenging.
Many works have focused on solving the aforementioned task under simpler problem settings. Unsupervised Pose-aware Part Decomposition (UPPD) [kawana] utilizes object shape annotations as a substitute for pose annotations. PartMobility [partMobility] utilizes multiple-frame point clouds of the same object under different joint states. However, these methods still face limitations when confronted with scenarios where shape information is unavailable or when dealing with single-frame data. To the best of our knowledge, Equi-Articulated-Pose (EAP) [PLSEE] is the only work that has tackled this task with single-frame point cloud as input and without shape or pose annotations on a synthetic dataset. EAP guides the network to learn part-by-part reconstruction of the input shapes by combining disentangled information, such as canonical part shapes, object structure, and part-level poses, in a self-supervised manner. To achieve such disentanglement, EAP extracts part-level $\mathrm{SE}(3)$-equivariant shape feature of a local region, instead of object-level $\mathrm{SE}(3)$-equivariant one, from an input and part-level poses. Since part-level poses are not given in nature, EAP requires iterative updates of such poses. It also uses an inner iterative operation, Slot-Attention [slot-attention], for segmenting parts. These iterative operations sacrifice inference speed.
@@
PartMobility [partMobility]
UPPD [kawana]
EAP [PLSEE]
@@ 2 Related Works
Category-level rigid object pose estimation: This task focuses on predicting an unknown rigid object’s pose from images. NOCS [NOCS] predicts the per-pixel coordinates in canonical space from RGB-D images. Several methods [Shapo, CenterSnap, ShapePrior] further employ CAD models from ShapeNet dataset [ShapeNet] to generate shape templates and use iterative closest point (ICP) [ICP] for matching the pose. Commonly used backbone for this task is 3D graph convolution network (3DGCN) [3DGCN] and PointNet++ [pointnet++]. These methods require expensive large-scale dataset annotation. Some approaches attempt to accomplish this task in a self-supervised manner. With the CAD model available, several methods [AAE, self6d] render the predicted pose with the CAD model as a synthetic image and compare it with the input image. Some methods focus on the multi-view RGB images provided cases [multiplePCPose, Multi-view]. Especially, $\mathrm{SE}(3)$-eSCOPE [SE3Rigid] achieved this task with single-view input and without pose annotations or CAD models. They use the $\mathrm{SE}(3)$-equivariant backbone, Equivariant Point Net (EPN) [EPN], to simultaneously conduct $\mathrm{SE}(3)$-invariant shape reconstruction as a reference frame, and predict the $\mathrm{SE}(3)$-equivariant pose transformation which sends input to the reconstruction.
@@
Category-level articulated object pose estimation: This task focuses on predicting part-level pose, part-level segmentation, and joint information for unknown objects within known categories. Previous methods [ANCSH, opd, generalize_kinematic, captra] try to solve this task with RGB-D image or video input by directly estimating the part-level pose. Some methods [asdf, cadex, neuralPart] transfer the task into a movable shape reconstruction task with neural implicit representation [OccNet, DeepSDF] and predict the pose indirectly. Some methods [active_articulation, ditto] parameterize the joint movement with active interaction with articulated objects. To reduce the segmentation cost, [semi_augmentation] uses semantic segmentation annotation and transfers it into part segmentation to conduct semi-supervised learning. However, similar to the rigid object pose estimation, the cost of the dataset annotation limits the application of these methods. To solve this task in a self-supervised manner, UPPD [kawana] utilizes the annotation of object shape instead of the annotation of object pose. Some methods [partMobility, multibody] used multi-view observation with the same object in different joint states to predict the joint movement. EAP [PLSEE] solved such a task with a single-frame point cloud input and without shape or pose annotation. EAP repeats the process of segmenting each part, reconstructing the per-part $\mathrm{SE}(3)$-invariant shape, and predicting the per-part pose multiple times to gain a refined pose estimation. However, directly segmenting parts for inputs with different poses and shapes is challenging, often resulting in poor accuracy, and the inference speed is unsuitable for real-time applications.
@@
Articulated Object Dataset: Synthetic datasets of articulated objects such as Shape2Motion, SAPIEN, and PartNet [shape2motion, SAPIEN, partnet] are commonly used in the articulated object pose estimation. Compared to RGB-D images captured from the real world, these datasets lack the consideration of complicated real-world environments. HOI4D [hoi4d] collects multiple articulated and rigid object mesh data and RGB-D images in human-object interaction. However, due to the mismatch between the depth and RGB channels, a non-negligible amount of noise is present in their ground-truth annotation of part segmentation based solely on the RGB channels.
@@ 3 Method
Figure 2: Pipeline of OP-Align. At the object-level phase, for the input point cloud $\mathbf{X}$, we use the E2PN [e2pn] backbone to predict and select object-level pose $\mathbf{R}_{\mathrm{o}},\mathbf{t}_{\mathrm{o}}$ from pose candidates, and generate the canonical reconstruction $\mathbf{Y}$ by adding a learnable parameter called category-common base shape $\mathbf{Y}_{\mathrm{base}}$. At the part-level phase, two PointNets [PointNet] with shared weights predict the part segmentation probability $\mathbf{W}_{\mathrm{x}},\mathbf{W}_{\mathrm{y}}$, joint states $\mathbf{a}_{\mathrm{x}},\mathbf{a}_{\mathrm{y}}$, joint pivots $\mathbf{c}_{\mathrm{x}},\mathbf{c}_{\mathrm{y}}$, and joint directions $\mathbf{d}_{\mathrm{x}},\mathbf{d}_{\mathrm{y}}$ for object-level aligned input $\mathbf{R}_{\mathrm{o}}\mathbf{X}+\mathbf{t}_{\mathrm{o}}$ and reconstruction $\mathbf{Y}$, to generate part-level alignment $\mathbf{R}_{\mathrm{d}},\mathbf{R}_{\mathrm{a}},\mathbf{T}_{\mathrm{a}}$ that aligns each part of $\mathbf{X}$ to the corresponding part of $\mathbf{Y}$ as part-level aligned inputs $\mathbf{Z}$.
Category-level articulated object pose estimation can be defined as follows. Given a point cloud $\mathbf{X}\in\mathbb{R}^{3\times N}$ of an articulated object consisting of $P$ parts, we assign each point to a part, predict the rotation and translation for each part, and provide the pivot and the direction for each joint. To solve this problem, our model predicts each point’s segmentation probability $\mathbf{W}\in\mathbb{R}^{P\times N}$, each joint’s pivot and direction $\mathbf{\{}\mathbf{c}_{[i]}\in\mathbb{R}^{3},\mathbf{d}_{[i]}\in\mathbb{R}^{3}\mid i\in\{1,2,\dots,J\}\}$, and the rotation and the translation for each part $\{\mathbf{R}_{[i]}\in\mathrm{SO}(3),\mathbf{t}_{[i]}\in\mathbb{R}^{3}\mid i\in\{1,2,\dots,P\}\}$. During training, we assume that the number of parts $P$ and the type of joints (revolute or prismatic) are given. Specifically, OP-Align assumes that each joint connects two independent parts, resulting in J=P1J=P-1 joints, which cover most of the articulated object categories found in daily environments.
The pipeline of OP-Align is shown in Fig. 2. At the object-level phase, OP-Align initially employs Efficient $\mathrm{SE}(3)$-equivariant Point Net (E2PN) [e2pn] for object-level pose selection from a discretization of the $\mathrm{SE}(3)$ group, and generate canonical reconstruction. At the part-level phase, two PointNets (PNs) [PointNet] with shared weights perform part segmentation and joint parameters estimation separately for the input aligned with object-level pose and the canonical reconstruction. The obtained joint parameters generate the part-level alignment between the input and the canonical reconstruction, aligning each part of the input with its corresponding part of the reconstruction by simulating the joint movement.
@@
Notice in this section, for a rank $n$ tensor $A$, we denote the $(i_{1},i_{2},\dots,i_{n})$-element (a rank 0-tensor) as $A_{[i_{1},i_{2},\dots,i_{n}]}$. Moreover, we use NumPy [numpy] like notation to extract a tensor from $A$ (but each index starts from $1$). For example, $A_{[i_{1}]}$ denote the $i_{1}$-th rank (n1)(n-1) tensor along the first axis and $A_{[:,i_{2}]}$ denote the $i_{2}$-th rank (n1)(n-1) tensor along the second axis.
@@ 3.1.1 Expansion from rigid objects to articulated objects
To solve the rigid object pose estimation in a self-supervised manner, $\mathrm{SE}(3)$-eSCOPE [SE3Rigid] utilizes a $\mathrm{SE}(3)$-equivariant backbone to disentangle shape and pose by generating an $\mathrm{SE}(3)$-invariant shape reconstruction and selecting $\mathrm{SE}(3)$-equivariant pose from candidates in a discretization of the $\mathrm{SE}(3)$ group for aligning the reconstruction and input. They observed that poses of $\mathrm{SE}(3)$-invariant reconstructions for objects in the same category are often consistently aligned. However, for articulated objects, each part’s pose is also influenced by joint movement. This complexity renders the reconstruction generated by the $\mathrm{SE}(3)$-eSCOPE unable to maintain consistent poses for all the parts.
@@ 3.1.2 Weighted Point Cloud Distances
We combine part segmentation probability with the point cloud distance between the part-level aligned inputs and the reconstruction to learn part segmentation and part alignment simultaneously. To achieve this, we use weighted point cloud distances, and later, part segmentation probability will sometimes be set as weights. A commonly used point cloud distance is the chamfer distance (CD), and we also employ the Density-awarded Chamfer Distance (DCD) [DCD]. Given two point clouds $\mathbf{P}$ and $\mathbf{Q}$, the single-directional weighted CD (L1) and DCD from $\mathbf{P}$ to $\mathbf{Q}$ with the weight $\mathbf{w}$ are defined as
[…]
@@ 3.2 Object-level phase
In the Object-level phase, OP-Align performs object-level pose selection, following a methodology similar to $\mathrm{SE}(3)$-eSCOPE [SE3Rigid], and generate canonical reconstruction.
By feeding the input $\mathbf{X}$, E2PN [e2pn] backbone initially outputs the $\mathrm{SE}(3)$-equivariant feature $\mathbf{F}_{\mathrm{eqv}}\in\mathbb{R}^{D\times 60}$. This feature is generated by $60$ anchors representing different poses of the object. Here, $60$ is the number of elements of the icosahedral rotation group, a discretization of the 33D rotation group $\mathrm{SO}(3)$. Then, we max pool $\mathbf{F}_{\mathrm{eqv}}$ among anchor dimension to obtain a $\mathrm{SE}(3)$-invariant feature $\mathbf{f}_{\mathrm{inv}}\in\mathbb{R}^{D}$. We use PoseHead, consisting of multi-layer perceptron (MLP), to output per-anchor rotation and translation $\{(\mathbf{R}_{[i]},\mathbf{t}_{[i]})=\textrm{PoseHead}(\mathbf{F}_{\mathrm{eqv}[:,i]})\mid i\in\{1,2,\dots,60\}\}$. To obtain the canonical reconstruction $\mathbf{Y}\in\mathbb{R}^{3\times N}$, we also use an MLP called ReconHead and a learnable parameter $\mathbf{Y}_{\mathrm{base}}$ which represents the category-common base shape and is of the same size as $\mathbf{Y}$. The canonical reconstruction $\mathbf{Y}$ is obtained by adding the output of ReconHead and $\mathbf{Y}_{\mathrm{base}}$; $\mathbf{Y}=\textrm{ReconHead}(\mathbf{f}_{\mathrm{inv}})+\mathbf{Y}_{\mathrm{base}}$.
@@ 3.3 Part-level phase
OP-Align uses two PNs [PointNet] with shared weights to process the object-level aligned input $\mathbf{R}_{\mathrm{o}}\mathbf{X}+\mathbf{t}_{\mathrm{o}}$ and the reconstruction $\mathbf{Y}$ separately. These two PNs output the segmentation probabilities $\mathbf{W}_{\mathrm{x}},\mathbf{W}_{\mathrm{y}}\in\mathbb{R}^{P\times N}$, joint pivots $\mathbf{c}_{\mathrm{x}},\mathbf{c}_{\mathrm{y}}\in\mathbb{R}^{(P-1)\times 3}$, joint directions $\mathbf{d}_{\mathrm{x}},\mathbf{d}_{\mathrm{y}}\in\mathbb{R}^{(P-1)\times 3}$ and per-part joint states $\mathbf{a}_{\mathrm{x}},\mathbf{a}_{\mathrm{y}}\in\mathbb{R}^{(P-1)\times 2}$ from each joint, where subscripts $\mathrm{x}$ and $\mathrm{y}$ indicate outputs from $\mathbf{R}_{\mathrm{o}}\mathbf{X}+\mathbf{t}_{\mathrm{o}}$ and $\mathbf{Y}$ respectively. Here, joint state $\mathbf{a}_{*}$ represents joint angles for revolute joints and translation lengths for prismatic joints, and the dimension of the second axis of $\mathbb{R}^{(P-1)\times 2}$ reflects the assumption that each joint connect two parts. We define the part-level aligned inputs $\mathbf{Z}_{[j,i]}$, $j=1,2,\dots,P-1$, $i=1,2$, obtained by a relative transformation that aligns the $i$-th part connected to the $j$-th joint of $\mathbf{R}_{\mathrm{o}}\mathbf{X}+\mathbf{t}_{\mathrm{o}}$ with the corresponding part of $\mathbf{Y}$ by